# AuctionNet: A Novel Benchmark for Decision-Making in Large-Scale Games

### v1

@@ Abstract
Decision-making in large-scale games is an essential research area in artificial intelligence (AI) with significant real-world impact. However, the limited access to realistic large-scale game environments has hindered research progress in this area. In this paper, we present AuctionNet, a benchmark for bid decision-making in large-scale ad auctions derived from a real-world online advertising platform. AuctionNet is composed of three parts: an ad auction environment, a pre-generated dataset based on the environment, and performance evaluations of several baseline bid decision-making algorithms. More specifically, the environment effectively replicates the integrity and complexity of real-world ad auctions through the interaction of several modules: the ad opportunity generation module employs deep generative models to bridge the gap between simulated and real-world data while mitigating the risk of sensitive data exposure; the bidding module implements diverse auto-bidding agents trained with different decision-making algorithms; and the auction module is anchored in the classic Generalized Second Price (GSP) auction but also allows for customization of auction mechanisms as needed. To facilitate research and provide insights into the game environment, we have also pre-generated a substantial dataset based on the environment. The dataset contains trajectories involving 48 diverse agents competing with each other, totaling over 500 million records and occupying 80GB of storage. Performance evaluations of baseline algorithms such as linear programming, reinforcement learning, and generative models for bid decision-making are also presented as part of AuctionNet. We note that AuctionNet is applicable not only to research on bid decision-making algorithms in ad auctions but also to the general area of decision-making in large-scale games. AuctionNet has been applied in the NeurIPS 2024 competition, providing nearly ten thousand accurate and fair algorithm evaluations for 1500 teams and inspiring more diverse and innovative solutions. Code [^1] Alibaba Group retains full ownership rights to this benchmark.: https://github.com/alimama-tech/AuctionNet.
@@ 1 Introduction
In this paper, we present AuctionNet, a benchmark for bid decision-making in large-scale ad auctions derived from a real-world online advertising platform. AuctionNet is composed of three parts: an ad auction environment, a pre-generated dataset based on the environment, and performance evaluations of a couple of baseline bid decision-making algorithms. More specifically, the environment effectively replicates the integrity and complexity of real-world ad auctions with the interaction of several modules: the ad opportunity generation module employs deep generative models to bridge the gap between simulated and real-world data while mitigating the risk of sensitive data exposure; the bidding module implements diverse auto-bidding agents trained with different decision-making algorithms; and the auction module is anchored in the classic and popular Generalized Second Price (GSP) [9, 23, 7] auction but also allows customization of auction mechanisms as needed. To facilitate research and provide insights into the game environment, we also pre-generated a substantial dataset based on the environment. The dataset contains trajectories with 48 diverse agents competing with each other, totaling over 500 million records and 80GB in size. Performance evaluations of baseline algorithms such as linear programming, reinforcement learning, and generative models for bid decision-making are also presented as a part of AuctionNet.
We note that AuctionNet is applicable not only to research on bid decision-making algorithms in ad auctions but also to the general area of decision-making in large-scale games. It can also benefit researchers in a broader range of areas such as reinforcement learning, generative models, operational research, and mechanism design.
@@ 3 Ad Auction Environment
The environment effectively replicates the integrity and complexity of real-world ad auctions through the interaction of several modules: the ad opportunity generation module, the bidding module, and the ad auction module. To better simulate large-scale auctions in reality, a substantial number of ad opportunities are fed into the environment and configured with dozens of bidding agents. These ad opportunities are generated using a deep generative model to reduce the gap between the simulation environment and reality while avoiding the risks of sensitive data exposure. The agents are equipped with diverse and sophisticated auto-bidding algorithms.
@@
3.1 Ad Opportunity Generation Module.
The target of the ad opportunity generation module is to generate diverse ad opportunities similar to real online advertising data. The core of this module is the deep generative model, as shown in Figure 2. We aimed to adopt the diffusion model to generate ad opportunity but encountered difficulties with the denoising operation, which can yield unreasonable outputs. Therefore, we followed the approach of the Latent Diffusion Model (LDM) [25] to generate ad opportunity. LDM adds noise and performs denoising in the latent space using a diffusion model, and then generates data from the latent space with an encoder and decoder. Specifically, LDM maps the ad opportunity feature $u$ to a latent vector $y$ with the encoder and reconstructs this feature with the decoder during training. For generation, LDM samples a random latent vector from a normal distribution and then generates an ad opportunity feature based on this vector. Let $U\subset\mathbb{R}^{d}$ be the space of ad opportunity feature data $(u_{1},u_{2},\cdots,u_{K})$, where $d$ is the dimension of the original data and $K$ is the number of ad opportunities. Let $Y\subset\mathbb{R}^{d^{\prime}}$ be the latent space ($d^{\prime}<d$). The encoder and decoder are represented as $g_{\phi}$ and $h_{\psi}$, respectively, where $\phi$ and $\psi$ are the parameters. The function of the encoder $g_{\phi}$ is to obtain a latent representation of the original data as $g_{\phi}(u_{k})=(\mu_{k},\sigma_{k})$, where $y_{k}\sim\mathcal{N}(\mu_{k},\sigma_{k}^{2})$ and $y_{k}\in Y$ is the latent representation. In practice, the reparameterization trick [20] is applied to ensure that this operation is differentiable during backpropagation.
@@
Different from the original idea of VAE, where the latent variable $y\in Y$ is sampled from $\mathcal{N}(0,1)$ in the generation process, LDM uses a diffusion model in the latent space to generate the latent variable. In general, the idea behind the diffusion model is to add Gaussian noise to the original data to obtain variables that follow $\mathcal{N}(0,1)$ and to denoise from $\mathcal{N}(0,1)$ for generation. Given a latent variable $y$, we denote its noisy version after $p$ iterations as $y_{p}$. The diffusion model includes a network to predict noise $\epsilon_{\theta}(y_{p},p)$, and the loss function can be represented as
@@
Several popular algorithms in the auto-bidding area have been implemented as baselines, including the PID Controller [36], Online LP [15], IQL [21], Behavior Cloning [30], and Decision Transformer [8]. This facilitates researchers who are interested in quickly starting up and evaluating these baselines in a unified environment.
3.3 Ad Auction Module.
The task of the ad auction module is to determine the winner and the winning price given all the bids from agents for ad opportunities. The costs for agents will vary depending on the different auction rules. The most commonly discussed auction rule is the Generalized Second-Price (GSP) Auction, which stipulates that the winner pays a cost slightly higher than the second-highest bid rather than the highest bid. The auction module internally supports several popular auction rules, including GSP, for the convenience of researchers. Additionally, researchers can design specific auction rules tailored to their purposes using the interface of the auction module.
@@ 4.2 Pre-Generated Dataset
The dataset includes 21 advertising episodes, each containing more than 500,000 ad opportunities and divided into 48 steps. Each opportunity includes the top 48 agents [^2] Real-world data show that 48 agents can ensure competitive pressure for auto-bidding agent training. with the highest bids. The dataset comprises over 500 million records, totaling 80 GB in size. Each record includes information such as the predicted value, bid, auction, and impression results, among other details. The specific data format and data samples of the dataset are included in Appendix C.
@@ 6 Applications
AuctionNet has been applied to the NeurIPS 2024 competition "Auto-bidding in Large-Scale Auctions" [1]. The competition addressed the critical issue of making high-frequency bid decision-making in uncertain and competitive environments and attracted more than 1,500 teams from around the world to participate, lasting for 3 months. The ad auction environment, dataset, and baseline bid decision-making algorithms used in the competition are derived from this benchmark. The ad auction environment provided nearly ten thousand evaluations for the competition, offering participants accurate and fair performance assessments. The dataset and baseline algorithms allowed participants to quickly start the task and stimulated their creativity, leading to more diverse and innovative solutions, thus driving technological development in this area.
@@ 8 Conclusion and Limitations
We present AuctionNet, a benchmark for bid decision-making in large-scale ad auctions derived from a real-world online advertising platform. AuctionNet consists of three components: an ad auction environment augmented with a verified deep generative model, a pre-generated dataset based on this environment, and performance evaluations of several baseline bid decision-making algorithms. The AuctionNet not only provides researchers with the opportunity to study auto-bidding algorithms in large-scale auctions, but also helps researchers and practitioners in game theory, reinforcement learning, generative models, operations optimization, and other fields to solve a wide range of decision-making research problems. Regarding limitations, while the generated data in the AuctionNet environment and the real-world data are similar in general, there are biases in some details, and the performance of the generative model can be improved.
@@ G.3 Bidding Module
Appendix H Details of Generative Model

### v2

@@ Abstract
Decision-making in large-scale games is an essential research area in artificial intelligence (AI) with significant real-world impact. However, the limited access to realistic large-scale game environments has hindered research progress in this area. In this paper, we present AuctionNet, a benchmark for bid decision-making in large-scale ad auctions derived from a real-world online advertising platform. AuctionNet is composed of three parts: an ad auction environment, a pre-generated dataset based on the environment, and performance evaluations of several baseline bid decision-making algorithms. More specifically, the environment effectively replicates the integrity and complexity of real-world ad auctions through the interaction of several modules: the ad opportunity generation module employs deep generative networks to bridge the gap between simulated and real-world data while mitigating the risk of sensitive data exposure; the bidding module implements diverse auto-bidding agents trained with different decision-making algorithms; and the auction module is anchored in the classic Generalized Second Price (GSP) auction but also allows for customization of auction mechanisms as needed. To facilitate research and provide insights into the environment, we have also pre-generated a substantial dataset based on the environment. The dataset contains 10 million ad opportunities, 48 diverse auto-bidding agents, and over 500 million auction records. Performance evaluations of baseline algorithms such as linear programming, reinforcement learning, and generative models for bid decision-making are also presented as a part of AuctionNet. AuctionNet has powered the NeurIPS 2024 Auto-Bidding in Large-Scale Auctions competition, providing competition environments for over 1,500 teams. We believe that AuctionNet is applicable not only to research on bid decision-making in ad auctions but also to the general area of decision-making in large-scale games. Code [^1] Alibaba Group retains full ownership rights to this benchmark.: https://github.com/alimama-tech/AuctionNet.
@@ 1 Introduction
In this paper, we present AuctionNet, a benchmark for bid decision-making in large-scale ad auctions derived from a real-world online advertising platform. AuctionNet is composed of three parts: an ad auction environment, a pre-generated dataset based on the environment, and performance evaluations of a couple of baseline bid decision-making algorithms. More specifically, the environment effectively replicates the integrity and complexity of real-world ad auctions with the interaction of several modules: the ad opportunity generation module employs deep generative networks to bridge the gap between simulated and real-world data while mitigating the risk of sensitive data exposure; the bidding module implements diverse auto-bidding agents trained with different decision-making algorithms; and the auction module is anchored in the classic and popular Generalized Second Price (GSP) [9, 23, 7] auction but also allows customization of auction mechanisms as needed. To facilitate research and provide insights into the game environment, we also pre-generated a substantial dataset based on the environment. The dataset contains 10 million ad opportunities, 48 diverse auto-bidding agents, and over 500 million auction records. Performance evaluations of baseline algorithms such as linear programming, reinforcement learning, and generative models for bid decision-making are also presented as a part of AuctionNet.
We believe that AuctionNet is applicable not only to research on bid decision-making algorithms in ad auctions but also to the general area of decision-making in large-scale games. It can also benefit researchers in a broader range of areas such as reinforcement learning, generative models, operational research, and mechanism design.
@@ 3 Ad Auction Environment
The environment effectively replicates the integrity and complexity of real-world ad auctions through the interaction of several modules: the ad opportunity generation module, the bidding module, and the auction module. To better simulate large-scale auctions in reality, a substantial number of ad opportunities are fed into the environment and configured with dozens of bidding agents. These ad opportunities are generated using deep generative networks to reduce the gap between the simulation environment and reality while avoiding the risks of sensitive data exposure. The agents are equipped with diverse and sophisticated auto-bidding algorithms.
@@
3.1 The Ad Opportunity Generation Module
The target of the ad opportunity generation module is to generate diverse ad opportunities similar to real online advertising data with deep generative networks, as shown in Figure 2. We aimed to adopt the diffusion model to generate ad opportunity but encountered difficulties with the denoising operation, which can yield unreasonable outputs. Therefore, we followed the approach of the Latent Diffusion Model (LDM) [25] to generate ad opportunity. LDM adds noise and performs denoising in the latent space using a diffusion model, and then generates data from the latent space with an encoder and decoder. Specifically, LDM maps the ad opportunity feature $u$ to a latent vector $y$ with the encoder and reconstructs this feature with the decoder during training. For generation, LDM samples a random latent vector from a normal distribution and then generates an ad opportunity feature based on this vector. Let $U\subset\mathbb{R}^{d}$ be the space of ad opportunity feature data $(u_{1},u_{2},\cdots,u_{K})$, where $d$ is the dimension of the original data and $K$ is the number of ad opportunities. Let $Y\subset\mathbb{R}^{d^{\prime}}$ be the latent space ($d^{\prime}<d$). The encoder and decoder are represented as $g_{\phi}$ and $h_{\psi}$, respectively, where $\phi$ and $\psi$ are the parameters. The function of the encoder $g_{\phi}$ is to obtain a latent representation of the original data as $g_{\phi}(u_{k})=(\mu_{k},\sigma_{k})$, where $y_{k}\sim\mathcal{N}(\mu_{k},\sigma_{k}^{2})$ and $y_{k}\in Y$ is the latent representation. In practice, the reparameterization trick [20] is applied to ensure that this operation is differentiable during backpropagation.
@@ 3.1 The Ad Opportunity Generation Module
Different from the original idea of VAE [20], where the latent variable $y\in Y$ is sampled from $\mathcal{N}(0,1)$ in the generation process, LDM uses a diffusion model in the latent space to generate the latent variable. In general, the idea behind the diffusion model is to add Gaussian noise to the original data to obtain variables that follow $\mathcal{N}(0,1)$ and to denoise from $\mathcal{N}(0,1)$ for generation. Given a latent variable $y$, we denote its noisy version after $p$ iterations as $y_{p}$. The diffusion model includes a network to predict noise $\epsilon_{\theta}(y_{p},p)$, and the loss function can be represented as
@@ 3.2 The Bidding Module
Several algorithms in the auto-bidding area have been implemented as baselines, including the PID Controller [36], Online LP [15], IQL [21], Behavior Cloning [30], and Decision Transformer [8]. This facilitates researchers who are interested in quickly starting up and evaluating these baselines in a unified environment.
3.3 The Auction Module
The task of the auction module is to determine the winner and the winning price given all the bids from agents for ad opportunities. The costs for agents will vary depending on the different auction rules. The most commonly discussed auction rule is the Generalized Second-Price (GSP) Auction, which stipulates that the winner pays a cost slightly higher than the second-highest bid rather than the highest bid. The auction module internally supports several popular auction rules, including GSP, for the convenience of researchers. Additionally, researchers can design specific auction rules tailored to their purposes using the interface of the auction module.
@@ 4.2 Pre-Generated Dataset
The dataset contains 10 million ad opportunities, including 21 advertising episodes. Each episode contains more than 500,000 ad opportunities, divided into 48 steps. Each opportunity includes the top 48 agents [^2] Real-world data show that 48 agents can ensure competitive pressure for auto-bidding agent training. with the highest bids. The dataset comprises over 500 million records, totaling 80 GB in size. Each record includes information such as the predicted value, bid, auction, and impression results, among other details. The specific data format and data samples of the dataset are included in Appendix C.
@@ 6 Applications
AuctionNet has powered the the NeurIPS 2024 competition "Auto-bidding in Large-Scale Auctions" [1]. The competition addressed the critical issue of making high-frequency bid decision-making in uncertain and competitive environments and attracted more than 1,500 teams from around the world to participate, lasting for 4 months. The ad auction environment, dataset, and baseline bid decision-making algorithms used in the competition are derived from this benchmark. The ad auction environment provided nearly ten thousand evaluations for the competition, offering participants accurate and fair performance assessments. The dataset and baseline algorithms allowed participants to quickly start the task and stimulated their creativity, leading to more diverse and innovative solutions, thus driving technological development in this area.
@@ 8 Conclusion and Limitations
We present AuctionNet, a benchmark for bid decision-making in large-scale ad auctions derived from a real-world online advertising platform. AuctionNet consists of three components: an ad auction environment augmented with verified deep generative networks, a pre-generated dataset based on this environment, and performance evaluations of several baseline bid decision-making algorithms. The AuctionNet not only provides researchers with the opportunity to study auto-bidding algorithms in large-scale auctions, but also helps researchers and practitioners in game theory, reinforcement learning, generative models, operations optimization, and other fields to solve a wide range of decision-making research problems. Regarding limitations, while the generated data in the AuctionNet environment and the real-world data are similar in general, there are biases in some details, and the performance of the generative model can be improved.
@@ G.3 Bidding Module
Appendix H Details of Deep Generative Networks